# Modular Soft Wearable Glove for Real-Time Gesture Recognition and Dynamic 3D Shape Reconstruction

### v1

@@ I INTRODUCTION
Human hands play a pivotal role in daily tasks and professional operations due to their high degrees of freedom (DOFs) that enable exceptional flexibility and precise control [1]. However, conventional rigid-body sensing technologies have inherent limitations in adaptability and responsiveness [2], making it difficult to capture the full range of hand configurations. Wearable gloves embedded with soft sensors have emerged as an effective solution for measuring the physical deformations and movements of the hand and fingers [3].
Recently, several strategies have been introduced to enhance the sensing capabilities of soft gloves. One approach leverages vision-based systems using cameras and markers to capture hand poses. While this method benefits from advanced computer vision algorithms and can achieve high accuracy, its effectiveness is limited by environmental factors and occlusions [4, 5]. Another strategy utilizes the intrinsic properties of human hands, such as surface electromyography (sEMG) [6, 7] and crease amplification [8] to infer hand gestures. These approaches offer robust sensing by leveraging physiological signals, but they usually require tight skin contact, and the model needs to be trained by patient-by-patient. Consequently, variations in how users wear the glove can significantly affect the sensor’s performance and effectiveness. A more direct approach employs strain sensors, which convert mechanical deformations into electrical signals. These sensors are capable of capturing hand movements without relying on the assistance of cameras or requiring a perfect fit with the hand. Such strain sensors are further categorized according to the sensing principle, including resistive [9, 10, 11, 12, 13, 14, 15, 16, 17], capacitive [18, 19, 20, 21, 22, 23] and triboelectric [24, 25] sensors.
@@ IV-B Results and Discussion
We employ the average distance (AD) [26] metric to evaluate hand movement reconstruction performance, which quantifies the discrepancy between the estimated and ground truth coordinates of visual markers:

### v2

@@ I INTRODUCTION
Human hands play a pivotal role in daily tasks and professional operations due to their high degrees of freedom (DOFs) that enable exceptional flexibility and precise control [1, 2]. However, conventional rigid-body sensing technologies have inherent limitations in adaptability and responsiveness [3], making it difficult to capture the full range of hand configurations. Wearable gloves embedded with soft sensors have emerged as an effective solution for measuring the physical deformations and movements of the hand and fingers [4].
Recently, several strategies have been introduced to enhance the sensing capabilities of soft gloves. One approach leverages vision-based systems using cameras and markers to capture hand poses. While this method benefits from advanced computer vision algorithms and can achieve high accuracy, its effectiveness is limited by environmental factors and occlusions [5, 6]. Another strategy utilizes the intrinsic properties of human hands, such as surface electromyography (sEMG) [7, 8] and crease amplification [9] to infer hand gestures. These approaches offer robust sensing by leveraging physiological signals, but they usually require tight skin contact, and the model needs to be trained by patient-by-patient. Consequently, variations in how users wear the glove can significantly affect the sensor’s performance and effectiveness. A more direct approach employs strain sensors, which convert mechanical deformations into electrical signals. These sensors are capable of capturing hand movements without relying on the assistance of cameras or requiring a perfect fit with the hand. Such strain sensors are further categorized according to the sensing principle, including resistive[10, 11, 12, 13, 14, 15, 16, 17, 18], capacitive [19, 20, 21, 22, 23, 24], optical [25, 26] and triboelectric [27, 28] sensors.
@@ IV-B Results and Discussion
We employ the average distance (AD) [29, 39] metric to evaluate hand movement reconstruction performance, which quantifies the discrepancy between the estimated and ground truth coordinates of visual markers: